OpenStudy (anonymous):

Lets us all try to solve this logic-physics question Usha swims a 90 m long pool. she covers 180m in just 1 min. by swimming from one end to other in a straight path find the average speed and the average velocity of usha

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Avg Speed = ( Total Distance ) / (Total Time) = (90+90)/60 = 3 m/s. Avg Velocity = ( Total Displacement ) / (Total Time) = (90 + (-90) ) / (60) = 0 . Since the start and the end point of Usha are the same, her displacement is nothing but the shortest distance between the start and the end points i.e, zero.But distance is the actual length of path covered. Hence, it is 90 + 90.

mathslover (mathslover):

for transparency : http://in.answers.yahoo.com/question/index?qid=20110524052020AAWpFBc

OpenStudy (anonymous):

thanks maths